Example of Vector space and subspace in which it's all properties must satisfied.


1
Expert's answer
2021-06-08T11:57:52-0400

Consider the vector space "\R^3" over field "\R" with usual addition

"(x_1,y_1,z_1)+(x_2,y_2,z_2)=(x_1+x_2,y_1+y_2,z_1+z_2)"

and multiplication by a scalar "\alpha\in\R:\ \alpha\cdot(x,y,z)=(\alpha x, \alpha y, \alpha z)."

Let us show that "W= \{(x,-3x,2x)|x\in\mathbb R\}" is a subspace of "\mathbb R^3". Let "\alpha, \beta\in\mathbb R,\ (x,-3x,2x), (y,-3y,2y)\in W." Then "\alpha (x,-3x,2x)+\beta (y,-3y,2y)= (\alpha x,-3\alpha x,2\alpha x)+(\beta y,-3\beta y,2\beta y)=
 (\alpha x+\beta y,-3\alpha x-3\beta y,2\alpha x+2\beta y)= (\alpha x+\beta y,-3(\alpha x+\beta y),2(\alpha x+\beta y))\in W."

Therefore, "W= \{(x,-3x,2x)|x\in\mathbb R\}" is a subspace of "\mathbb R^3".